Find the maximum number of points of intersection of 7 straight lines and 5 circle when 3 straight lines are parallel and 2 circles are concentric

Grade:11

1 Answers

Arun
6 years ago
Dear Raghav
 
L & L intersection = 7C2 - 3C2 = 18
L & C intersection
2 concentric circles = 14 *2 = 28
3rd Circle = 14 + (2*2) = 18
4th Circle = 14 +(3*2) = 20
5th circle = 14 + (4*2) = 22
Thus
Point of intersections = 18 +28 +18 + 20+ 22 = 106
Hwnce final answer is 106.
 
Regards,
Arun ( AskIITians Forum Expert)